OpenStudy (anonymous):

Solve the equation. Show your work. 16^3=4^x

OpenStudy (anonymous):

16 = 4^2

OpenStudy (anonymous):

so 4^(2*3) = 4^x 4^6 = 4^x x = 6

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

hope...this helps..))

OpenStudy (anonymous):

Yup...Why,....u have answer key...))

OpenStudy (anonymous):

sure do lol jk

OpenStudy (anonymous):

hm...hm....u check.....

OpenStudy (anonymous):

im just playing yahoo

OpenStudy (anonymous):

i am ...correct..lol

OpenStudy (anonymous):

lol :)